# An Investigation of the Current Star Formation Rate of Star-forming Galaxies at z < 0.5 with FADO

### v1

@@ 2.3 Spectral fitting
The PyQSOFit is a dedicated data package for fitting SDSS spectra, primarily used for fitting quasar spectra and calculating the central supermassive black hole masses in quasars. It can also be applied to galaxy spectra, such as for calculating the SFR of galaxies. Based on PyQSOFit, Fu (2021) improved it to become qsofitmore, which is no longer limited to fitting SDSS spectra but can be used to analyze spectral data files obtained from other devices, such as LAMOST. qsofitmore has shown good performance in spectral fitting and analysis and has become increasingly popular in quasar and galaxy spectral studies (Fu et al., 2022; Jin et al., 2023; Ding et al., 2023; Chen et al., 2024). This improved Python package not only fits the emission line information of target galaxy spectra but also uses Monte Carlo (MC) method to obtain the flux errors of the fitted components. It should be noted that qsofitmore is based on the BC03 model (Bruzual & Charlot, 2003) for spectral fitting, which does not simultaneously consider the radiations of stars and nebulae to decompose the spectral components, as FADO does. This allows us to compare the SFR obtained considering nebulae radiation in FADO with that obtained by qsofitmore to demonstrate whether there is a significant difference in higher redshifts.Additionally, in qsofitmore, we take into account the potential contamination of galaxy spectra by the host galaxy. Thus, we remove the contribution from the host galaxy to obtain smoother galaxy spectra. As shown in the appendix.
FADO is considered a self-consistent and dynamically calculated spectral fitting tool that takes into account the contribution of nebular radiation to emission line luminosities. FADO is the first population spectral synthesis(pss) code employing genetic differential evolution optimization. By identifying the main nebular features in galaxies through the observed Balmer line luminosities and equivalent widths (EWs), as well as the shape of the continuum around the Balmer and Paschen jump regions, FADO realizes the estimation of nebular emission lines. This is achieved through the full-consistency mode (FC), Nebular-Continuum mode (NC), and STellar mode (ST) (Gomes & Papaderos, 2017). The main difference between FADO and other pure stellar codes (such as STARLIGHT and qsofitmore) is that FADO considers the contribution of nebular emission lines in galaxies rather than solely estimating the SFR of the target galaxy using stellar radiation. The advantage of FADO lies in its ability to simultaneously fit and estimate multiple parameters, such as age, metallicity, stellar mass, and dust content in galaxies. This allows for a more comprehensive analysis of the physical properties of galaxies. Both qsofitmore and FADO use the MC method to assign spectral flux errors, and their basic template spectra are derived from the BC03 model. Additionally, they can both perform extinction correction, dereddening, and K-correction on spectra under given celestial coordinates and redshifts, ensuring comparability with the fitting of qsofitmore and maintaining control and uniformity in the basic spectral processing conditions.
@@
Therefore, we conducted fittings on the spectra of 2575 galaxies obtained from Section 2.1 using FADO and qsofitmore, respectively, and conducted preliminary inspection and screening of the fitting results (e.g., fitted fluxes of H and H). In regions of active star formation, there are sufficiently young and hot stars that provide enough energy for the ionization of hydrogen. These ionized gases capture photons and undergo recombination transitions, leading to simultaneous emission of H and H. Additionally, the significant electromagnetic radiation emitted during star formation is heavily attenuated by dust obscuration. Since dust absorption and attenuation are more efficient at longer wavelengths than shorter ones, this results in a decrease in the H-H ratio (Calzetti et al., 2000). Theoretically, an H-H ratio of 2.86 is commonly regarded as a typical intrinsic extinction indicator reflecting the dust absorption effects on hydrogen emission lines (Hummer & Storey, 1987; Gaskell & Ferland, 1984). Therefore, we corrected for intrinsic extinction using an H-H ratio of 2.86 for all our data when calculating SFR (Cardelli et al., 1989). Both FADO and qsofitmore software come with Milky Way extinction corrections, thus intrinsic extinction correction for galaxies is necessary. The detailed procedure of the extinction correction is presented in the Appendix.
Table 1 shows five criteria were used to select galaxies suitable for subsequent analysis: (1) z < 0.5; (2) high spectral quality and appropriate S/N; (3) selection based on the BPT diagram; (4) successful fitting of H; and (5) successful fitting of both main emission lines, H and H. Criterion (4) was employed to directly compare the differences in H fitting between the two spectral fittings without sacrificing too many samples. The distribution of this data is quantified in the Left panel of Figure 3. Criterion (5) was implemented to compare the differences in SFR between the two fittings. Before calculating SFR, intrinsic extinction correction is necessary to ensure the reliability of the estimated SFR, hence both H and H needed to be successfully fitted. The distribution of this data is quantified in the Right panel of Figure 3. The corresponding selected samples consisted of 2575 and 2215 spectra, respectively. After excluding three outliers, we obtained a golden sample containing 2212 galaxy spectra, all of which met the requirements of this investigation. The complete spectral selection criteria and the resulting samples are presented in Table 1. For the estimation of SFR, the number of galaxies in the golden sample was 1807. The reason for sample loss was the intrinsic extinction correction, and the specific data are shown in Figure 8.A comparison of the emission line fittings between FADO and qsofitmore in criteria (4) and (5) reveals that FADO performed better, fitting more emission lines and reducing sample loss. It is worth noting that the last column in the table represents the number of samples that satisfied the corresponding criteria and were successfully fitted by both FADO and qsofitmore, which are the data truly used for sample selection in this study. During the selection process, the sharp decrease and scarcity of samples were greatly influenced by the fitting results of qsofitmore, as it imposes stricter criteria on spectral fittings. This indicates that compared to the recently developed FADO spectral fitting method, previous methods for fitting galaxy spectra have certain limitations under the assumption of reasonable and excellent spectral fitting capabilities of FADO. After obtaining the golden sample, necessary parameter comparative analyses were conducted, and the results are presented in the next section.
| Criteria | FADO | qsofitmore | Combined | |
SDSS-DR18 z 0.5 |
1,730,947 | 1,730,947 | 1,730,947 |
S/N above red line and z 5 |
36,109 | 36,109 | 36,109 |
BPT diagram selection |
2,575 | 2,575 | 2,575 |
H fitted |
2,575 | 2,212 | 2,212 |
H and H fitted Final gold sample |
2,563 | 1,807 | 1,807 |
The distribution of data points in this survey with respect to redshift is illustrated in Figure 3. The left panel represents the data after BPT diagram selection, while the right panel depicts the subset that satisfies both the BPT diagram selection and the fitting of H6563 lines. It is evident from both histograms that there is a declining trend, with a notable difference in the number of data points between the redshift intervals of 0-0.1 and 0.4-0.5. Despite the careful selection of an appropriate number of data points in each redshift interval in Figure 1, the discrepancy in the data is more pronounced in Figure 3. This is attributed to the fact that the redshift interval of 0-0.1 corresponds to nearby galaxies, which are typically well observed with advanced techniques and thorough research methods, resulting in higher-quality spectra that are less likely to be filtered out by FADO and qsofitmore. Conversely, the redshift interval of 0.4-0.5 exceeds the median redshift of the SDSS survey (around 0.1) (Abazajian et al., 2009), leading to a obviously drop in the data shown in Figure 3. Moreover, the requirement of simultaneously fitting [N II] 6583, [O III] 5007, H6563, and H4861 lines in the BPT diagram poses a stringent constraint, particularly for the data in the 0.4-0.5 redshift range, especially for H6563. Around a redshift of 0.5, H6563 may fall beyond the maximum wavelength coverage of the SDSS, resulting in its failure to be fitted, thereby further reducing the data. The decrease in data in the right panel is due to the stringent fitting criteria of qsofitmore and FADO, which may erroneously fit flux values for some low-quality spectra, leading to data reduction. This reduction is particularly evident in the 0.4-0.5 redshift interval in the right panel, where most of the data are removed due to the fitting of H6563 and H4861 lines. Consequently, the main focus of this investigation is on the sample concentrated between 0 and 0.4, as the data in the 0.4-0.5 interval are sparse. However, there are better data trends in the following figures, which will be shown in next section.
3 results and analysis
The purpose of this investigation is to compare the SFR derived from qsofitmore and FADO, aiming to evaluate whether FADO can provide a more reasonable SFR by excluding nebular contamination. As mentioned in Section 2.2, H6563 is currently considered the most suitable estimator for galaxy SFR. Therefore, this investigation utilizes the H6563 luminosity to calculate the SFR of the golden sample. Equation 4 and Equation 5 are employed to compute the optimal SFR values.
@@
Where is the luminosity distance, the best fitted flux of after extinction of H6563, the luminosity of 6563 and , the multiplication factor given by (Miranda et al., 2023).Calibrated to (Chabrier, 2003) IMF. Among them, qsofitmore and FADO use consistent calibration and consistent calculation methods to calculate SFR.
The estimation results indicate that FADO yields SFR values for the golden sample, which are statistically nearly half of those provided by qsofitmore. As illustrated in Figure 4, 950 galaxies (45%) exhibit lower SFR values from FADO compared to those from qsofitmore. The black dashed line in the figure represents equality between the SFR values obtained from both methods, with error bars derived using Monte Carlo methods. The top panel comprises five subplots quantifying the comparison of fits across different redshift intervals, where the black dashed line signifies equality between the two fitting tools. Top panel of Figure 4 shows a good trend of changes with redshift. It can be seen that from z = 0-0.1 to z = 0.3-0.4, as the redshift increases, the black dotted line have more and more data under, which shows that more and more data points indicate that H6563 fitted by FADO is smaller than H6563.This aligns with our expectation since qsofitmore, in computing SFR, does not account for the nebular component, thus overestimating the emission intensity from star-forming activities, leading to inflated SFR estimates. Figure 5 quantifies all data points below the diagonal line, with the redshift range of 0.2-0.4 exhibiting notable differences, with approximately 69% of the data falling below the qsofitmore fit. Figure 5 preliminarily demonstrates the sensitivity of the SFR derived from FADO fits to the H6563 line in the range of z = 0.2 0.4, consistent with the inference by Miranda et al. (2023). However, it is crucial to note that our sample remains too limited, particularly in the redshift range of 0.4-0.5. This limitation arises partly due to the selected redshift range, preventing us from obtaining a sufficient number of spectroscopic parent samples solely from SDSS optical spectra for fitting. Additionally, the suboptimal spectral observation quality of galaxies within the chosen redshift range by SDSS leads to the exclusion of the majority based on criterion (4) (see Section 2.1). Thus, more high-quality spectra are needed to be collected to expand the golden sample, to better elucidate the trends depicted in Figure 5. Furthermore, during the spectral fitting stage, we observed that qsofitmore imposes stringent conditions on spectral fitting, with the majority of spectra failing to exhibit the emission lines that should ideally be present, possibly due to the SED fitting model used as a reference by qsofitmore (Fu, 2021). The high requirements for spectral quality by qsofitmore also contribute significantly to the scarcity of the golden sample. Additionally, we conducted a KS test on the x and y coordinates of Figure 4 to assess whether they follow the same distribution. The KS statistic: 0.178, P-value: 2.01E-32, indicates a significantly small P-value, suggesting significant distributional differences between the two datasets. The Ha emissions fitted by qsofitmore and FADO do not conform to the same distribution.
To further investigate the trend of the ratio of H6563 fluxes from FADO and qsofitmore with redshift, we present in Figure 6 the relationship between the ratio of H6563 fluxes and redshift. Since the calculation of SFR in this study is based on the H6563 emission line luminosity, to avoid reducing our parent sample due to the joint fitting of H6563 and H4861, we adopt the H6563 / H4861 ratio to quantitatively characterize the extent of bias introduced by the nebular component to SFR estimation in past studies. The relationship between this ratio and redshift can reveal how the contribution of nebular radiation to galaxy emission changes with increasing redshift. After linearly fitting the sample, we found a slight decrease in the vertical axis with increasing redshift. This may suggest that at higher redshifts, the effect of nebulae on SFR estimation becomes more pronounced, consistent with the inference of (Miranda et al., 2023). However, due to the narrow range of selected redshifts and the small sample size, this trend is not apparent. Therefore, more high-quality spectra of intermediate to high-redshift galaxies are needed in the future to amplify this trend and verify the universality of this phenomenon. The large error in the data in Figure 6 may be attributed to the significant estimation error resulting from FADO’s use of a Monte Carlo code combined with many models, which will be analyzed in the next section. It can be observed from Figure 6 that regardless of whether errors are included, the final slope is negative, indicating a decrease in the vertical axis with increasing horizontal axis. The decrease in the vertical axis implies that is overall smaller than , which is consistent with our initial speculation. This is because FADO removes some radiation originally from nebulae when fitting the H6563 emission line, fitting only the radiation from stars, resulting in a more accurate and smaller flux.
Next, we attempt to quantify the differences between the two fitting methods in computing SFR. They employ the same formula for SFR calculation, as shown in Equation 4 and Equation 5, thus ensuring comparability of the SFR calculation data. The intrinsic extinction correction of galaxies has been considered in the SFR calculation, and detailed methods are provided in the Appendix. The SFMS (star-forming main sequence) is plotted against stellar mass, which is estimated by FADO. There exists a close relationship between stellar mass and SFR, with current research indicating a positive linear logarithm relationship. Since SFR calculations are based on extinction-corrected data, both and H4861 need to be fitted simultaneously by both methods. This requires a re-evaluation of the data, and Figure 7 depicts the scenario where both H6563 and H4861 are fitted. It can be observed that there are very few data points remaining in the redshift range of 0.4-0.5, which can be almost neglected. This is due to the stringent fitting criteria of FADO and qsofitmore, coupled with the poor spectral quality in the redshift range of 0.4-0.5, which is insufficient to meet the requirements of this study. The left panel of Figure 8 illustrates the SFMS diagram from FADO, while the right panel shows the SFMS diagram from qsofitmore. The two lines in Figure 8 represent the slopes of the SFR calculated by FADO and qsofitmore, respectively. To facilitate comparison of their differences, these two slopes are juxtaposed. The data from this survey also adhere to the SFMS relationship established by other researchers, where the overall SFR increases with increasing stellar mass. However, we observe that the amplitude of SFR calculated by FADO increases much more rapidly with increasing stellar mass compared to that calculated by qsofitmore. The SFR calculated by FADO tends to align more closely with a proportionate upward trend, whereas that by qsofitmore tends to approach a straight line. This indicates that the SFR calculated by FADO exhibits a better trend, capturing a flux of H6563 that conforms more closely to theoretical models of the SFMS trend, further underscoring the robustness of FADO. The right panel of Figure 8 displays the SFR calculated by qsofitmore, indicating almost no upward trend, with data distribution tending towards a flat line and lacking a discernible linear trend, contrary to the SFMS prediction by FADO.
Finally, we conclude that approximately 45 of the data points indicate that the H6563 emission line fitted by FADO is smaller than that fitted by qsofitmore, which preliminarily suggests that FADO successfully subtracts some nebular emissions. Additionally, by examining the trend of the ratio between FADO fitting results and qsofitmore fitting results with redshift, we tentatively infer that nebular contamination increases with higher redshifts. Therefore, it is anticipated that when studying the SFR at high redshifts, the influence of nebular emissions should be removed.

### v3

@@ 2.3 Spectral fitting
The qsofitmore is a Python module dedicated to fitting and analyzing spectra of quasars and galaxies (Fu, 2021). The module has shown good performance in spectral fitting and analysis and has become increasingly popular in quasar and galaxy spectral studies (Fu et al., 2022; Jin et al., 2023; Ding et al., 2023; Chen et al., 2024). The qsofitmore not only fits the emission lines of the spectra but also uses the Monte Carlo (MC) method to obtain the uncertainties of the fitted components. It should be noted that the module is based on the BC03 model (Bruzual & Charlot, 2003) for spectral fitting, which does not simultaneously consider the stellar and nebular radiations to decompose the spectral components, as FADO does. This allows us to compare the obtained by FADO and qsofitmore to demonstrate whether a significant difference exists.
FADO is a self-consistent spectral fitting tool that can dynamically calculate the parameters of a galaxy spectrum. It estimates the contribution of nebular radiation from emission line luminosity. FADO estimates nebular emission lines through the Full-Consistency, Nebular-Continuum, and STellar mode (Gomes & Papaderos, 2017). The main difference between FADO and other pure stellar codes (such as STARLIGHT and qsofitmore) is the consideration of nebular emission or not. Both qsofitmore and FADO adopt basic template spectra derived from the BC03 model and estimate parameter uncertainties with the MC process. In addition, both of them can perform extinction correction, de-redshift, and K-correction on spectra under given coordinates and redshifts. These may ensure the comparability of spectral fitting. A comparison of the spectral fitting by qsofitmore and FADO for the same galaxy is displayed in Appendix B.
Following the comparison above between the two fitting tools, we conducted spectral fitting on the 6558 galaxies obtained in Section 2.1 using qsofitmore and FADO, respectively. To consider intrinsic extinction correction, the two fittings must give physically meaningful results of the flux of and , as well as their corresponding uncertainties. Then 6,546 and 6,523 galaxies were filtered out by qsofitmore and FADO, respectively. Finally, a total of 6,511 galaxies were selected as the Golden Sample, in which the spectra of galaxies are successfully fitted by both FADO and qsofitmore. Their statistics of the redshift and quantity are displayed in Figure 3. Notably, there is an obvious scarcity of quantity in the redshift interval from 0.3 to 0.4. This is due to the deficiency of the SDSS survey on the higher-redshift galaxy. The redshift range of 0.3 0.4 exceeds the median redshift of the SDSS survey (around 0.1, Abazajian et al., 2009), leading to a noticeable drop in the sample quantity. Besides, may fall beyond the maximum wavelength coverage of the SDSS around a redshift of 0.4, resulting in its failure to be fitted. Therefore, the requirement of the BPT diagnosis constraints more strictly on sample volume, particularly in the redshift range of 0.3 0.4. Although galaxies in this range are relatively sparse, these samples are important and cannot be ignored in this work.
2.4 Intrinsic extinction correction
@@
There are sufficient young and hot stars in the star-forming regions that provide enough energy for the ionization of hydrogen. The ionized gas captures photons and undergoes recombination transitions, leading to simultaneous emission of and . However, cold gas and dust heavily attenuate the significant electromagnetic radiation emitted during star formation. Since attenuation of cold matter is more efficient at shorter wavelengths than longer ones, this results in a decrease in the - ratio (Calzetti et al., 2000). Commonly, a - ratio of 2.86 is regarded as a typical intrinsic extinction corrector (Gaskell & Ferland, 1984; Hummer & Storey, 1987; Cardelli et al., 1989). Therefore, the intrinsic extinction of Golden Sample was corrected with the ratio.
3 analysis and results
This section compares the flux of Golden Sample fitted by both qsofitmore and FADO to assess whether the two fittings are consistent. The nebular contribution in the wavelength range of is calculated to examine the variation with redshift. The SFR is also estimated with the flux fitted by FADO to evaluate whether FADO can provide a more reasonable SFR by excluding nebular contamination.
3.1 Comparison of flux
As is shown in Figure 4, the flux obtained from FADO is statistically similar to that from qsofitmore, which is consistent with the result in Figure 2 of Miranda et al. (2023). KS test was performed on the qsofitmore, and FADO fitted flux to assess whether they follow the same distribution. The KS statistic is 0.068, and the p-value is , suggesting statistical uniformity between the two fitted subsets. Further, the flux was binned by redshift with the interval of 0.1 to explore whether there exists a tendency with redshift between the two fittings. The results in Figure 5 indicate no clear hint of difference caused by redshift up to at least z 0.4. In general, the flux fitted by both qsofitmore and FADO do not exhibit noticeable differences, which is consistent with the results of Miranda et al. (2023). The Middle Difference is .
3.2 Nebular Contribution
To further analyze and quantify the possible nebular contribution, the fitted nebular spectra were extracted from the FADO fitting outputs. The fitted nebula components were binned by redshift and averaged within the corresponding redshift intervals. These averaged nebular spectra displayed in panel (a) of Figure 6 increase with the increase of redshift at the centre wavelength of (black dashed line), except for the averaged spectra in the interval of 0.3 < z < 0.4 (green line). The decline of the green averaged spectra relative to the red one would be due to the lack of the sample and the low quality of the spectra. We expect the green averaged spectra to be raised after the expansion of the higher-redshift sample. The shape of the nebula radiation in Figure 6 panel (a) is all uniform, it comes from the FADO fitting template. For details, refer to Gomes & Papaderos (2017).
Figure 6 panel (b) shows that the nebula ratio of the fitted nebular spectra was also integrated within the wavelength range from 6,400 Å to 6,800 Å to quantify the impact on . The Nebula ratio is equal to the Nebula Model fitted in FADO divided by the Stellar Model. This method visualizes the percentage of nebulae in the entire spectrum. For a detailed explanation, please see Appendix B. Figure 6 panel (b) presents the linear analytic expressions corresponding to the four redshift intervals, plotted in the colours of their respective scatter points, along with an overall trend represented by a black solid line. Calculating the cumulative nebular contribution at wavelengths between 6,400 Å and 6,800 Å for the samples in different redshift intervals reveals a proportional increase from to . This indicates that for the samples used in this work, the nebular contribution increases, in fact, with the redshift, directly confirming the hypothesis proposed by Miranda et al. (2023).
However, the increase in redshift is not stable and has a large dispersion. The Pearson correlation coefficient can be used to better quantify this issue. The Pearson correlation coefficient for is -0.077, for is 0.189, for is 0.013, and for it is -0.065. The total Pearson correlation coefficient for the four redshift bins is 0.181. Overall, the nebula ratio has a more obvious trend with redshift, with a slope of 1.46. The reason for the difference in the sensitivity of the nebula ratio to redshift in the above four redshift bins may be the instability of gases (especially the density of hydrogen) at medium and low redshifts.
The hydrogen density in galaxies is also related to the epoch of the galaxy. Early-type galaxies contain far more gas than late-type galaxies (Rhee et al., 2013). Since this work uses the H emission line to compare the differences between two fitting codes and utilizes the H flux to estimate the SFR, the hydrogen density value has a significant impact on the nebula ratio. Figure 9 of Rhee et al. (2013) summarizes the effect of redshift on hydrogen density. The H2 density at is defined by (Zwaan et al., 2005), at by (Martin et al., 2010), at by (Delhaize et al., 2013), around by (Freudling et al., 2010), and around by (Rhee et al., 2013). Therefore, the H2 density at should be the lowest, which corresponds to the trend in panel (b) Figure 6, where the nebula ratio decreases with redshift in . At , due to the depletion of cold gas, the galaxy evolution becomes static, which leads to a decreasing trend in the nebula ratio.
For galaxies with stellar masses greater than , the content of H2 + He follows a trend of first increasing, then stabilizing, and eventually decreasing in the redshift range , as shown by (Geach et al., 2011; Pavesi et al., 2018). So, we would expect that the relationship between the nebular ratio and the redshift would yield more significant results if the stellar mass and galaxy type were consistent.
3.3 Stellar Mass Effect
For the nebula ratio, there is still an invisible variable, the stellar mass of the galaxy, which affects the redshift relationship. This is due to the limitations of human sky photography techniques. This effect can be obtained from the z-SFMS diagram.
Despite its limitations, the flux is still widely considered the most accurate estimator of galaxy star formation rate for nearby galaxies compared to other estimators, and it is the primary tracer used in this analysis. Therefore, this investigation utilizes the luminosity to calculate the SFR of the golden sample. Equation (3) and Equation (4) are employed to compute the optimal SFR values.
@@ 3.3 Stellar Mass Effect
Where is the luminosity distance, the best fitted flux of after extinction of , the luminosity of and , the multiplication factor given by (Miranda et al., 2023).Calibrated to Chabrier (2003) IMF. Among them, qsofitmore and FADO use consistent calibration and consistent calculation methods to calculate SFR.
Next, we attempt to quantify the differences in SFR calculations using FADO. Since qsofitmore does not have the capability to calculate stellar mass, only FADO data are used to plot the SFMS. The SFR is calculated using the equations as shown in Equation (3) and Equation (4). Intrinsic extinction corrections for galaxies have been considered during the SFR calculations. A strong correlation between stellar mass and SFR is observed, with current studies indicating a positive log-linear relationship between them. As the SFR calculation is based on extinction-corrected data, reasonable values for and must be obtained from the fits. The left panel of Figure 7 shows the SFMS diagram derived from FADO, while the right panel shows the variation of the SFMS diagram with redshift. The red and black lines in Figure 7 represent the slopes of the SFR calculated from the left and right panels, respectively. As they are derived from the same diagram, the slopes are consistent. The data from this investigation also conform to the established SFMS relationship from other studies, in which overall SFR increases with stellar mass, consistent with the findings of Miranda et al. (2023) in Figure 9. Ultimately, Figure 7 shows that the relationship between stellar mass and SFR is described as .
In the right panel of Figure 7, it is evident that stellar mass and redshift exhibit a strong positive correlation, as does SFR with redshift. This clear positive relationship arises from the early universe (high redshift), where gas density and energy levels were higher, allowing gas to collapse more efficiently to form stars. In the later stages of the universe (mid-to-low redshift), as gas is depleted—excluding external factors such as gas replenishment or galaxy mergers—star formation rates decline accordingly. This decline begins around the epoch known as ”cosmic noon” (z 2) (Behroozi et al., 2013). Across the redshift range 0-8, a positive linear relationship is observed between and SFR (Madau & Dickinson, 2014; Behroozi et al., 2013). Stellar formation density also increases with redshift in the ranges 0-0.8 and 0-5 (Hopkins & Beacom, 2006), and stellar mass shows a strong positive correlation with redshift in the range 0-4 (Ilbert et al., 2013).
Redshift | Mass Median | SFR Median |
0–0.1 | ||
0.1–0.2 | ||
0.2–0.3 | ||
0.3–0.4 |
Notes. We present the median value and the standard deviation of the stellar mass and SFR for the four redshift intervals.
Table 1 quantifies the changes in stellar mass and SFR with redshift, as shown in the right panel of Figure 6. As in previous sections, the redshift is divided into four intervals to observe the corresponding changes in stellar mass and SFR. As predicted, for the data from this investigation, both stellar mass and SFR increase steadily with redshift. It should be noted that observational biases may affect the z-SFMS relationship. At higher redshifts, galaxies that can be detected tend to be brighter, while fainter galaxies are harder to observe, leading to the possible exclusion of low-mass, high-redshift galaxies. The brighter galaxies are associated with a higher number of active O-type and B-type stars, which leads to a higher SFR. Consequently, galaxies at higher redshifts tend to have higher stellar masses and higher SFRs. To mitigate the effects of these factors on the interpretation of the relationship between redshift and the nebula ratio, the Nebula Ratio is normalized by the stellar mass to balance these influences. One way to mitigate these observational biases is to match the luminosities of high-redshift galaxies with those of various mass galaxies at lower redshifts, thus achieving a controlled comparison.
Figure 8 offsets the effect of stellar mass on the nebula ratio. Compared with Figure 6 panel (b), the overall trend does not seem to change significantly, and the trend of the nebula ratio with increasing redshift weakens. However, the dispersion is still large, which may be related to the distribution of gas in space.
3.4 Redshift VS Nebula Ratio
Figure 9 quantifies the impact of redshift on the average nebula ratio for the current SFG sample while also considering the influence of stellar mass. The vertical axis has been normalized. The yellow data points represent the case where the effect of stellar mass is not taken into account, while the pink data points indicate the scenario where the stellar mass effect is considered. The corresponding analytical expressions are presented in Equation (5).
Equation (5) displays the standard quadratic function fitting expression. The quadratic function fitting provides the best results based on the tests conducted. For the orange trend line, which does not consider the stellar mass, the coefficients are , , and . For the pink trend line, which accounts for the stellar mass, the coefficients are , , and . By comparing the values of these two datasets, it is observed that the value for the data points without the consideration of stellar mass is 0.950, whereas the value for the data points with the stellar mass effect is 0.877. Therefore, it is concluded that the impact of stellar mass on the nebula ratio is relatively small.